A person conducts a contest in which the listeners phone jn the answer to the contest question. If the person estimates that the probability of a caller giving the correct answer is 0.28,

a) What is the probability that the third caller is the winner?
b) What is the expected number of calls before a correct answer is received?

Answer by Boreal(15235)

a. This would be 0.72*0.72*0.28 or p(1-p)^2
=0.1452
=
b. E(X)=1/p=3.57
It is a geometric distribution but the pdf is x(1-x)^(n-1), where n=3, so that is where the square term comes from.